OpenStudy (anonymous):

2y = -x + 5 is the equation of a line that passes through the point (1, 2) and has a slope of -1/2. (True of false)

OpenStudy (mertsj):

What is the slope?

OpenStudy (anonymous):

i got x/4 after multiplying the given set

OpenStudy (mertsj):

Solve the equation for y

OpenStudy (mertsj):

Have you heard that before?

OpenStudy (anonymous):

but it tells me the slope is -1/2 so i must be wrong

OpenStudy (mertsj):

Did you solve the equation for y?

OpenStudy (anonymous):

yes, y=-1/2+5/2

OpenStudy (mertsj):

Where did the x go?

OpenStudy (anonymous):

-1/2x my bad..

OpenStudy (mertsj):

The coefficient of x is the slope so yes, we see that the slope is - 1/2

OpenStudy (anonymous):

so we know that part is true..

OpenStudy (mertsj):

Now you also want to know if the point (1,2) is on the line. So replace x with 1 and y with 2 and see if it makes the sentence true.

OpenStudy (mertsj):

I guess I am all mixed up. I thought the equation was 2y=-x+5

OpenStudy (mertsj):

Where did you get 4y=x+5?

OpenStudy (anonymous):

4y=-1x+5 4 does equal -1 + 5 so would this be true although i excused x and y?

OpenStudy (mertsj):

Where did you get 4y?

OpenStudy (anonymous):

multiplying 2 by 2y?

OpenStudy (mertsj):

Ohhhhhhl, I see. You don't know how to substitute into an equation.

OpenStudy (anonymous):

guess not.

OpenStudy (mertsj):

Did you every play on a sports team where you had a coach?

OpenStudy (mertsj):

ever

OpenStudy (anonymous):

no im handicapped

OpenStudy (mertsj):

Do you every watch sports on tv and see a coach substitute one player for another?

OpenStudy (anonymous):

yes.

OpenStudy (mertsj):

So one player comes out and another replaces him. That is what happens when you substitute 2 for y and 1 for x. y comes out and 2 replaces it. x comes out and 1 replaces it so we have:

OpenStudy (mertsj):